OpenStudy (anonymous):

The length of a rectangle is equal to three times its width. If the perimeter is equal to 200 feet, what is the length of the rectangle?

OpenStudy (anonymous):

 The perimeter is the sum of all sides. You're going to want to add the 4 sides together. Since there are two of each side, you're going to have to count them all. 2(3w)+2w=200 (I multiplied them by two to count in all 4 sides.) 6w+2w=200 8w=200 (solve for x by dividing both sides by 8) w=25 The width is 25 feet, and the length is 75 feet (three times the width)